if the temperature of the helium balloon were increased from 30 celcius to 35 celcius and the volume of the ballon only expanded from 0.47 L to 1.55 L, how much pressure would be inside the ballon?
1
Expert's answer
2020-04-20T15:06:20-0400


Let us assume that the initial pressure be p1 and final pressure be p2 so lets find out a relation between them

As the balloon is a closed body so air will escape through it which means that the number of moles initially and finally will be same

PV=nRT

n=PV/RT

as this n will be same on both sides so lets equate

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"

"\frac{P_1}{P_2}=\frac{1.55\times303}{0.47\times308}=3.24"